An electron is accelerated to a speed of    3 x106 m/s in an unknown potential V. What is the magnitude of the electric potential V?
Answer #1

here

V = W / q0

V = 3 * 10^6 * 9.1 * 10^-31 / ( 1.6 * 10^-19)

V = 1.7 * 10^-5 volts